one of the factors of 6x^2+7x-10 is

Answer by checkley71(8403)

6X^2+7X-10=0
(6X-5)(X+2)=0
6X-5=0
6X=5
X=5/6 ONE ANSWER.
X=2=0
X=-2 ANOTHER ANSWER.